OpenStudy (anonymous):

Justin drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took hours. When Justin drove home, there was no traffic and the trip only took hours. If his average rate was miles per hour faster on the trip home, how far away does Justin live from the mountains?

OpenStudy (anonymous):

trip there took 9 hours, trip back took 4 hours. average rate was 40 mph faster on trip home, how far away does Justin live from the mountains?

OpenStudy (anonymous):

so the real question is distance....how about d d=4 hours for one trip and d=9 for the other trip: d can't equal 9 and 4 so need more info maybe need s for speed d=4s and d=9s 4s can't equal 9s that makes not sense but we know the shorter trip was s+40 so d=(4(s+40) and d=9s If we set those equal to each other I'll bet it can be solved 9s=4(s+40)